OpenStudy (anonymous):

How do I solve: x+y-6 x+2y=9 Please show the steps involved. Thank you.

myininaya (myininaya):

Multiply 1st equation by -1 (first i think you are missing an equation sign there) then add the equations together

myininaya (myininaya):

x+y=6 => -x-y=-6 +(x+2y=9) ------- 0+y=3 y=3

myininaya (myininaya):

y=3 and x+y=6 => x+3=6 => x=3

myininaya (myininaya):

is that right x+y=6 or x+y=-6?

OpenStudy (anonymous):

subtract 1st eqn fro 2nd eqn. Then u get y=3 When applying the value of y in the 1st eqn x+3=6 x=3

OpenStudy (anonymous):

Can someone explain to me how -y + 2y = y?

myininaya (myininaya):

-1+2=1

myininaya (myininaya):

2-1=1

OpenStudy (anonymous):

Ah, I see, thank you. Also, why did you change the first equation to all negative? Is there any extra steps I need to take in order to do that or do I just change it all to negative naturally?

myininaya (myininaya):

i just wanted one of the variables to cancel when i added the equations together

myininaya (myininaya):

so i seen the x's would cancel if one of them were negative

myininaya (myininaya):

so if I multiply one term of an equation by -1 then I multiply all the terms of that equation by -1

OpenStudy (anonymous):

Okay, thank you very much :)